# BoxFusion: Reconstruction-Free Open-Vocabulary 3D Object Detection via Real-Time Multi-View Box Fusion

### v2

@@ Abstract
Open-vocabulary 3D object detection has gained significant interest due to its critical applications in autonomous driving and embodied AI. Existing detection methods, whether offline or online, typically rely on dense point cloud reconstruction, which imposes substantial computational overhead and memory constraints, hindering real-time deployment in downstream tasks. To address this, we propose a novel reconstruction-free online framework tailored for memory-efficient and real-time 3D detection. Specifically, given streaming posed RGB-D video input, we leverage Cubify Anything as a pre-trained visual foundation model (VFM) for single-view 3D object detection by bounding boxes, coupled with CLIP to capture open-vocabulary semantics of detected objects. To fuse all detected bounding boxes across different views into a unified one, we employ an association module for correspondences of multi-views and an optimization module to fuse the 3D bounding boxes of the same instance predicted in multi-views. The association module utilizes 3D Non-Maximum Suppression (NMS) and a box correspondence matching module, while the optimization module uses an IoU-guided efficient random optimization technique based on particle filtering to enforce multi-view consistency of the 3D bounding boxes while minimizing computational complexity. Extensive experiments on ScanNetV2 and CA-1M datasets demonstrate that our method achieves state-of-the-art performance among online methods. Benefiting from this novel reconstruction-free paradigm for 3D object detection, our method exhibits great generalization abilities in various scenarios. The code will be released at project page.
@@ 1 Introduction
Recently, embodied intelligence has attracted great interest, with many well-developed applications, such as embodied navigation [2, 42, 46], mobile manipulation [45, 39]. These embodied applications are mainly based on scene understanding (e.g., 3D object detection [28, 29], semantic segmentation [9, 22], and depth estimation [40, 17]). Current methods all reconstruct the scene first for scene understanding with dense representations, such as point clouds[24], voxels[47], etc. However, we found that a sparse representation of reality, which only preserves the position, scale, and semantics of objects’ bounding boxes, as a scene representation, is sufficient to complete the downstream tasks, such as scene graph generation [12] and collision detection [33].
@@
Primary research in 3D object detection has focused predominantly on representations of 3D point clouds or meshes in the past. As inputs to these methods, point clouds inevitably undergo sparsification operations for computational feasibility, which introduce detection deficiency caused by lower resolution. Learning-based methods aim to incorporate inductive biases into models to mitigate this issue [24, 28, 5]. Yet, the substantial discrepancy between 3D representations and physical reality restricts their applicability primarily to large, prominent objects. Therefore, some point-based methods [25] have attempted to use images as auxiliary inputs to enhance the performance. Recently, with the development of visual foundation models (VFMs) [27, 10, 19], researchers have started to leverage VFMs to incorporate rich semantic information from 2D images into point cloud-based methods, enabling open-vocabulary detections  [1, 7, 21]. While these methods are still dependent on dense point cloud reconstruction, which is prone to being affected by noise (e.g., accumulated errors in the reconstruction or noise in the sensors) or storage limitations of large-scale scenarios. Furthermore, these methods typically follow a two-stage paradigm that requires complete dense point cloud reconstruction as inputs, which significantly hinder real-time applications. To address these challenges, some recent works have explored online perception methods. These methods either introduce a memory adapter [38] to store the reconstructed point cloud or utilize 2D foundational models to extract fine-grained object masks and semantics from each view and then fuse them with the reconstructed dense point cloud for open-vocabulary perception [37, 31]. However, these online perception methods still rely on dense point-cloud reconstruction as input, which significantly increases computational overhead and degrades real-time performance, making it difficult to meet the requirements of real-time interaction in many downstream tasks.
@@
To overcome these challenges, in this work, we propose a novel reconstruction-free online framework specifically designed for memory-efficient and real-time open-vocabulary 3D object detection. Our approach is inspired by many embodied tasks, such as navigation and manipulation, which do not require a complete reconstruction of the scene but rather focus on detecting and understanding the objects within it. As shown in Figure 1, by eliminating the need for dense point-cloud reconstruction, we can achieve faster and more efficient 3D object detection while maintaining high accuracy. Specifically, for streaming RGB-D video input, we leverage Cubify Anything [14] as a pre-trained visual foundation model (VFM) for efficient single-view 3D object proposal generation. These proposals are enriched with open-vocabulary semantics by using CLIP [27] to extract the semantic feature of the detected instance. An association module then aggregates multi-view proposals corresponding to the same instance through 3D Non-Maximum Suppression (NMS) and 2D box correspondence matching. Since the proposals from Cubify Anything are local and view-specific, it is of great necessity to aggregate and optimize the associated proposals from different views of the same object. Therefore, to further enhance multi-view consistency, we employ an optimization module, an IoU-guided efficient random optimization technique, which is grounded by particle filtering to overcome the highly non-linear optimization of object boxes. In this way, we can enforce multi-view consistency of the proposals of 3D bounding boxes while minimizing computational complexity to achieve real-time 3D object detection.
@@
Extensive experiments on the CA-1M and ScanNetV2 datasets demonstrate that our method achieves state-of-the-art performance among online open-vocabulary 3D detection methods. Benefiting from its reconstruction-free paradigm, our approach exhibits exceptional generalization capabilities and enables real-time performance even in large-scale environments exceeding 1000 square meters, paving the way for the deployment in large-scale and challenging scenarios. In summary, our contributions are as follows:
@@
Figure 2: Method Overview of BoxFusion. (a) Given online RGB-D images with camera poses, we use Cubify Anything to generate bounding box proposals for each keyframe, and back-project them to image planes to get the image crops to obtain the open-vocabulary semantics with CLIP. (b) We employ an association module to perform 3D NMS for global boxes with proposals from the new keyframe, removing redundant boxes and associating those belonging to the same object. (c) With the associated candidate boxes of an object, we adopt random optimization based on particle filtering to fuse these candidate boxes into a multi-view consistent one using the IoU of the convex hulls of projected box corners. (e.g., there are three boxes for the highlight desk at three timestamps that only cover a part of it, and the boxes are not accurate enough). In this way, our method can efficiently detect fine-grained objects in real time without reconstruction.
@@ 2 Related Work
Traditional 3D object detection: Pioneered by PointNet [23], 3D object detection methods used to rely on dense labeled data corresponding to closed-set categories of objects. Point-based methods [24, 36, 20] follow the feature extraction techniques used in 2D object detection, and extract permutation-invariant point cloud features by 3D backbones, coupled with 3D detection heads. The follow-up works [6, 3, 13] boost the performance by incorporating graph neural networks, better proposal generation, and contextual information. backbones like Voxel-based methods [47, 28, 29], akin to 2D detection, convert point clouds into regular grids, equipped with 3D convolutional neural networks (CNNs) for feature extraction. Favorable detection on notable room-level objects (e.g., chairs and tables), but poor generalization on objects of unseen categories probably hinders the applications on more complicated tasks.
@@
Open-vocabulary 3D object detection: With the development of visual foundation model (VFM) like SAM [10] or DINO [19], recent works [21, 35, 7] first establish the correspondence between 3D data like point clouds and 2D RGB images, and subsequently employ vision-language model (VLM) like CLIP [30] to segment images for more fruitful object proposals and corresponding semantics. These methods demonstrate improvements in detecting objects of novel categories in 3D scenes and make cross-modal retrieval between images and text descriptions feasible. However, these methods still rely on 3D data like point clouds as the core representation, which often suffers from the relatively poor resolution of point clouds and limits the flexibility with high computational overheads. SpatialLM [32] encodes the reconstructed point cloud and detects all the objects queried by a designed text prompt. Another paradigm is to leverage the 2D image features for 3D object detection, predicting the 3D bounding box directly. Cubify Anything [14] predicts metric fine-grained 3D bounding boxes by using a novel large-scale dataset and Encoder-Decoder vision transformer (ViT). DetAny3D [43] proposes a novel 3D detection framework that directly predicts 3D bounding boxes for a single image, leveraging the rich image features from VFM. However, these methods are primarily designed for single-view 3D object detection and exhibit limited generalization capabilities in dynamic online scanning scenarios.
@@
Online 3D perception: Online 3D perception is a crucial task in robotics, enabling real-time scene understanding towards the surrounding environments.  [38] employs a memory-based adapter for online 3D perception, including semantic and instance segmentation and 3D object detection. However, it is limited to closed-set categories and does not support open-vocabulary detection. Recent works [41, 37, 31] exploit the potential of 2D VFM like SAM [10] or CropFormer [26] for online 3D instance segmentation via per-frame mask merging. However, multi-view fusion relying on 3D data representations (e.g., point clouds) inherently introduces sparsification artifacts and sensor-induced noise during reconstruction (e.g., depth holes from invalid measurements). Consequently, methods dependent on reconstructed point clouds to determine the spatial overlap of different masks are fragile, especially for small objects. Moreover, maintaining a dense 3D structure increases the computational overhead, and methods that rely on reconstructed point clouds to determine the spatial overlap of different masks are fragile, especially for small objects. Maintaining a dense 3D structure increases the computational overhead and reduces the running and memory efficiency in terms of online settings and larger-scale scenarios. Compared to these methods, our method is reconstruction-free and more efficient in diverse scenarios.
@@ 3 Methodology
In this section, we first introduce the overall pipeline of our method. The insight of our method is to leverage foundation models to generate single-view fine-grained 3D bounding boxes and fuse them into multi-view consistent global 3D bounding boxes. In this way, the proposed reconstruction-free pipeline is able to generate scene-level 3D bounding boxes of fine-grained objects.
@@ 3.1 Overview
As shown in Figure 2, our method consists of three main components: single-view prediction, box association, and box fusion. Given RGB-D posed images as inputs, we first predict single-view 3D bounding box proposals using Cubify Anything [14] for each keyframe, and transform the boxes from camera to world coordinates. These 3D proposals are then projected back to 2D in order to crop the images and use CLIP [27] for semantic features. To associate the candidate boxes of the same instance, we first use 3D NMS to filter spatially overlapping boxes, coupled with a 2D correspondence matching module to associate small objects, since proposals of small objects are more likely to be adjacent but not close enough to have overlaps. Finally, we fuse these proposals into a single global 3D bounding box for each object in the scene. Specifically, inspired by [44], we employ the IoU-guided random optimization equipped with a pre-sampled particle swarm templates (PST) for better online efficiency and robustness to the highly-nonlinear optimization of box fusion. By this means, the proposed method is real-time and efficient for fine-grained scene-level 3D bounding box generation.
@@ 3.2 Proposal Generation
In the era of foundation models, 3D object detection has been significantly boosted by leveraging large-scale pre-trained models. In this work, we adopt Cubify Anything [14] for fine-grained box proposal generation. Cubify Anything is a state-of-the-art method that can generate 3D bounding boxes from a single RGB-D image, and it is trained on a large-scale dataset with exhaustive objects.
To be specific, given the successive RGB-D inputs $\{(I_{t},D_{t})\}_{t=1}^{N}$ with $N$ images in total, where $I_{t}$ is the RGB frame and $D_{t}$ is the depth frame, we first obtain $M$ 3D bounding box proposals $\{\bar{B}^{i}_{t}\}_{i=1}^{M}$ for each image using Cubify Anything. Each box $\bar{B}^{i}_{t}$ is represented in the camera coordinate system. To transform the boxes to the world coordinate system, we apply the camera pose transformation $\mathcal{T}_{t}$ to each box $\bar{B}^{i}_{t}$, and obtain the global 3D boxes $\{\mathcal{G}^{i}_{t}\}_{i=1}^{M}$. In order to obtain the boxes with the corresponding semantics, we project the 3D boxes back to the 2D image planes $I_{t}$ and crop the corresponding image regions according to the projected box corners. The cropped images are then fed into a pre-trained CLIP model $\Theta$ [27] to obtain the semantic features $\{F_{i}\}_{i=1}^{N}$ of each box proposal as shown in Eq. 1.
$F^{i}_{t}=\Theta(P(I_{t},K_{t},\{\bar{B}^{i}_{t}\})),$ (1)
where $K_{t}$ denotes the camera intrinsics and $P$ indicates the projection and image cropping operation.
@@ 3.3 Association Module
In this section, we introduce the box association module to associate the 3D bounding boxes of the same instance across different views. The association is crucial for maintaining consistent global 3D bounding boxes since numerous bounding boxes are generated from each single view. The spatial association and correspondence association are introduced as follows.
Figure 3: Visualization of multi-view box fusion using random optimization with the IoU of the intersecting projected convex hulls between the global box and single-view proposals. Three single-view proposals of a cabinet (bottom) are fused into a global one (top).
@@ 3.3.1 Spatial Association
Previous methods [37, 31] mainly exploit spatial overlap for association, which is valid for most room-level common objects like chairs and desks, but can not generalize to small objects (e.g., a TV remote on the table). The reason is that proposals for most small objects generated from different methods are usually adjacent but not spatially overlapped, which may result in the ignorance of detection for small objects. To address this challenge, we propose a novel two-stage association module: 3D Non-Maximum Suppression (NMS) for associating spatially overlapping boxes, and a 2D correspondence matching module for associating the remaining boxes, which are designed to associate fine-grained objects.
Concretely, given the local box proposals $B^{i}_{t}$, we add these to global bounding boxes $\mathcal{G}_{t-1}$ and form current global boxes $\mathcal{G}_{t}$ at the frame $t$. We need to determine whether any of the global boxes correspond to the other and remove the redundant ones. Compared to the axis-aligned NMS used in previous methods [24, 20], we aim to generate scene-level oriented 3D bounding boxes. Therefore, we adopt the NMS for oriented bounding boxes. To be specific, we first sort $\mathcal{G}_{t}$ by the corresponding scores from Section 3.2. For each box $\mathcal{G}^{i}_{t}$ indexed by the descending order of scores, we calculate the IoU between ${G}^{i}_{t}$ and the other boxes in $\mathcal{G}_{t}$. The IoU is calculated based on the convex hull of the 3D bounding boxes, which is more suitable for oriented boxes.
For online efficiency, we uniformly sample $O_{n}$ points in the convex hull of each pair of 3D bounding boxes, and only those that are likely to have overlap are considered by checking if the points on the edges are included in each other. For a pair of 3D bounding boxes $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$ ($i>j$ in the descending order of scores), the 3D convex-based IoU $\Omega_{ij}$ is defined as Eq. 2. We manage a candidate list $\Psi$ for all global boxes $\mathcal{G}_{t}$. We filter $\mathcal{B}_{j}$ if $\Omega_{ij}$ is higher than the threshold $\tau_{3d}$ and $\mathcal{B}_{j}$ is then associated with $\mathcal{B}_{i}$ if the camera direction is larger than $\tau_{r}$ or the camera translation is larger than $\tau_{t}$ compared to all candidate boxes in $\Psi_{i}$.
$\Omega_{ij}=\frac{\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{B}_{i}\cup\mathcal{B}_{j})}{\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{B}_{i})+\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{B}_{j})},$ (2)
where $\mathcal{C}$ is the indicator function that returns 1 if the point $O_{k}\in\mathcal{B}_{i}$, otherwise 0 and $\mathcal{B}_{i}\cup\mathcal{B}_{j}$ denotes the union of convex hulls of 3D bounding box $\mathcal{B}_{i}$ and $\mathcal{B}_{j}$.

### v3

@@ Abstract
Open-vocabulary 3D object detection has gained significant interest due to its critical applications in autonomous driving and embodied AI. Existing detection methods, whether offline or online, typically rely on dense point cloud reconstruction, which imposes substantial computational overhead and memory constraints, hindering real-time deployment in downstream tasks. To address this, we propose a novel reconstruction-free online framework tailored for memory-efficient and real-time 3D detection. Specifically, given streaming posed RGB-D video input, we leverage Cubify Anything as a pre-trained visual foundation model (VFM) for single-view 3D object detection, coupled with CLIP to capture open-vocabulary semantics of detected objects. To fuse all detected bounding boxes across different views into a unified one, we employ an association module for correspondences of multi-views and an optimization module to fuse the 3D bounding boxes of the same instance. The association module utilizes 3D Non-Maximum Suppression (NMS) and a box correspondence matching module. The optimization module uses an IoU-guided efficient random optimization technique based on particle filtering to enforce multi-view consistency of the 3D bounding boxes while minimizing computational complexity. Extensive experiments on CA-1M and ScanNetV2 datasets demonstrate that our method achieves state-of-the-art performance among online methods. Benefiting from this novel reconstruction-free paradigm for 3D object detection, our method exhibits great generalization abilities in various scenarios, enabling real-time perception even in environments exceeding 1000 square meters. The code will be released at project page.
@@ 1 Introduction
Recently, embodied intelligence has attracted great interest, with many well-developed applications, such as embodied navigation [3, 49, 54], mobile manipulation [53, 46]. These embodied applications are mainly based on scene understanding (e.g., 3D object detection [34, 35], semantic segmentation [13, 28], pose estimation [10] and depth estimation [47, 23]). Current methods reconstruct the scene first for scene understanding with explicit representations, such as point clouds[30], voxels[55], etc. However, inspired by [22], we found that a sparse representation of reality, which only preserves the position, scale, and semantics of objects’ bounding boxes, as a scene representation, is sufficient to support the downstream embodied tasks, such as scene graph generation [16] and collision detection [40].
@@
Primary research in 3D object detection has focused predominantly on representations of 3D point clouds or meshes in the past. As inputs to these methods, point clouds inevitably undergo sparsification operations for computational feasibility, which introduce detection deficiency caused by lower resolution. Learning-based methods aim to incorporate inductive biases into models to mitigate this issue [30, 34, 8]. Yet, the substantial discrepancy between 3D representations and physical reality restricts their applicability primarily to large, prominent objects. Therefore, some point-based methods [31, 2, 4] have attempted to use images as auxiliary inputs to enhance the performance. Recently, with the development of visual foundation models (VFMs) [33, 14, 25], researchers have started to leverage VFMs to incorporate rich semantic information from 2D images into point cloud-based methods, enabling open-vocabulary detections  [1, 11, 27]. While these methods are still dependent on dense point cloud reconstruction, which is prone to being affected by noise (e.g., accumulated errors in the reconstruction or noise in the sensors), or storage limitations of large-scale scenarios.
Furthermore, these methods typically follow a two-stage paradigm that requires complete point cloud reconstruction as inputs, which significantly hinder real-time applications. To address these challenges, some recent works have explored online perception methods. These methods either introduce a memory adapter [45] to store the reconstructed point cloud or utilize 2D foundational models to extract fine-grained object masks and semantics from each view and then fuse them with the reconstructed dense point cloud for open-vocabulary perception [44, 38]. However, these online perception methods still rely on dense point-cloud reconstruction as input, which significantly increases computational overhead and degrades real-time performance, making it difficult to meet the requirements of real-time interaction in many downstream tasks.
@@
To overcome these challenges, in this work, we propose a novel reconstruction-free online framework specifically designed for memory-efficient and real-time open-vocabulary 3D object detection. Our approach is inspired by many embodied tasks, such as navigation and manipulation, which do not require a complete reconstruction of the scene but rather focus on detecting and understanding the objects within it. As shown in Figure 1, by eliminating the need for dense point-cloud reconstruction, we can achieve faster and more efficient 3D object detection while maintaining high accuracy in large-scale scenarios. Specifically, for streaming RGB-D video input, we leverage Cubify Anything [19] as a pre-trained visual foundation model (VFM) for single-view 3D object proposal generation. These proposals are enriched with open-vocabulary semantics by using CLIP [33] to extract the semantic features of the detected objects. An association module then aggregates multi-view proposals corresponding to the same instance through 3D Non-Maximum Suppression (NMS) and 2D box correspondence matching. Since the proposals from Cubify Anything are local and view-specific, it is of great necessity to aggregate and optimize the associated proposals from different views of the same object. Therefore, to further enhance multi-view consistency, we employ an optimization module, an IoU-guided efficient random optimization technique, which is grounded by particle filtering to overcome the highly non-linear optimization of object boxes. In this way, we can enforce multi-view consistency of the proposals of 3D bounding boxes while minimizing computational complexity to achieve real-time 3D object detection.
@@
Extensive experiments on the CA-1M and ScanNetV2 datasets demonstrate that our method achieves state-of-the-art performance among online open-vocabulary 3D detection methods. Benefiting from the reconstruction-free paradigm, our approach exhibits exceptional generalization capabilities and enables real-time performance even in large-scale environments exceeding 1000 square meters, paving the way for the deployment in large-scale and challenging scenarios. In summary, our contributions are as follows:
@@
Figure 2: Method Overview of BoxFusion. (a) Given online RGB-D images with camera poses, we use Cubify Anything to generate bounding box proposals for each keyframe, and project them to image planes to get the image crops and obtain open-vocabulary semantics with CLIP. (b) We employ an association module to perform 3D NMS for global boxes with proposals from the new keyframe, removing redundant boxes and associating those belonging to the same object. (c) With the associated candidate boxes of an object, we adopt random optimization based on particle filtering to fuse these candidate boxes into a multi-view consistent one using the IoU of the convex hulls of projected box corners. (e.g., the highlight desk is partially covered by three bounding boxes at three timestamps, which are not accurate enough). In this way, our method can efficiently detect fine-grained objects in real time without reconstruction.
@@ 2 Related Work
Traditional 3D object detection: Pioneered by PointNet [29], 3D object detection methods used to rely on dense labeled data corresponding to closed-set categories of objects. Point-based methods [30, 43, 26] follow the feature extraction techniques used in 2D object detection, and extract permutation-invariant point cloud features by 3D backbones, coupled with 3D detection heads. The follow-up works [9, 5, 17] boost the performance by incorporating graph neural networks, better proposal generation, and contextual information. Voxel-based methods [55, 34, 35], akin to 2D detection, convert point clouds into regular grids, equipped with 3D convolutional neural networks (CNNs) for feature extraction. However, these methods focus on notable room-level objects (e.g., chairs and tables) and show poor generalization on objects of unseen categories, hindering the applications on more complicated tasks.
@@
Open-vocabulary 3D object detection: With the development of visual foundation model (VFM) like SAM [14] or DINO [25], recent works [27, 2, 42, 11] first establish the correspondence between 3D data like point clouds and 2D RGB images, and subsequently employ vision-language model (VLM) like CLIP [36] to segment images for more fruitful object proposals and corresponding semantics. These methods demonstrate improvements in detecting objects of novel categories in 3D scenes and make cross-modal retrieval between images and text descriptions feasible. However, these methods still rely on 3D data like point clouds as the core representation, suffering from the relatively poor resolution and noise of point clouds. This results in high computational overheads and limits the flexibility. SpatialLM [39] encodes the reconstructed point cloud and detects all the objects queried by a designed text prompt. Another paradigm is to leverage the 2D image features for 3D object detection, predicting the 3D bounding box directly. Cubify Anything [19] predicts metric fine-grained 3D bounding boxes by using a novel large-scale dataset and Encoder-Decoder vision transformer (ViT). DetAny3D [50] proposes a novel 3D detection framework that directly predicts 3D bounding boxes for a single image, leveraging the rich image features from VFM. However, these methods are primarily designed for single-view 3D object detection and exhibit limited generalization capabilities in scene-level online scanning scenarios.
@@
Online 3D perception: Online 3D perception is a crucial task in robotics, enabling real-time scene understanding towards the surrounding environments.  [45] employs a memory-based adapter for online 3D perception, including semantic and instance segmentation and 3D object detection. However, it is limited to closed-set categories and does not support open-vocabulary detection. Recent works [48, 44, 38] exploit the potential of 2D VFM like SAM [14] or CropFormer [32] for online 3D instance segmentation via per-frame mask merging. However, multi-view fusion relying on 3D data representations (e.g., point clouds) inherently introduces sparsification artifacts and sensor-induced noise during reconstruction (e.g., depth holes from invalid measurements). Consequently, these methods, which utilize the spatial overlap of different masks for aggregation, are fragile, especially for small objects. The explicit reconstruction also increases the computational overhead and reduces the running and memory efficiency in terms of online settings and larger-scale scenarios. Compared to these methods, our method is reconstruction-free and more efficient in diverse scenarios.
@@ 3 Methodology
In this section, we first introduce the overall pipeline of our method. The insight of our method is to leverage foundation models to generate single-view fine-grained 3D bounding boxes and fuse them into multi-view consistent global 3D bounding boxes. In this way, the proposed reconstruction-free pipeline can efficiently generate scene-level 3D bounding boxes of fine-grained objects in real time.
@@ 3.1 Overview
As shown in Figure 2, our method consists of three main components: proposal generation, box association, and box fusion. Given RGB-D posed images as inputs, we first predict single-view and local 3D box proposals using Cubify Anything [19] for each keyframe, and transform the boxes from camera to world coordinates. These 3D proposals are then projected to 2D in order to crop the images and use CLIP [33] for semantic features. To associate the candidate boxes of the same instance, we first use 3D NMS to filter spatially overlapping boxes, coupled with a 2D correspondence matching module to associate small objects, since proposals of small objects are more likely to be adjacent but not close enough to have overlaps. Finally, we fuse these proposals into a single global 3D bounding box for each object in the scene. Specifically, inspired by [51, 52], we employ the IoU-guided random optimization equipped with a pre-sampled particle swarm templates (PST) for better online efficiency and robustness to the highly-nonlinear optimization of box fusion. In this way, the proposed method is robust and efficient for online open-vocabulary 3D object detection.
@@ 3.2 Proposal Generation
In the era of foundation models, 3D object detection has been significantly boosted by leveraging large-scale pre-trained models. In this work, we adopt Cubify Anything [19] for fine-grained box proposal generation. Cubify Anything is a state-of-the-art method that can generate 3D bounding boxes from a single RGB-D image, and it is trained on a large-scale dataset with exhaustive objects that are labeled with ground-truth boxes.
To be specific, given the successive RGB-D inputs $\{(I_{t},D_{t})\}_{t=1}^{N}$ with $N$ images in total, where $I_{t}$ is the RGB frame and $D_{t}$ is the depth frame, we first obtain $M$ 3D bounding box proposals $\{{B}^{i}_{t}\}_{i=1}^{M}$ for each image using Cubify Anything. Each box ${B}^{i}_{t}$ is represented in the camera coordinate system. To transform the boxes to the world coordinate system, we apply the camera pose transformation $\mathcal{T}_{t}$ to each box ${B}^{i}_{t}$, and obtain the global 3D boxes $\{\mathcal{G}^{i}_{t}\}_{i=1}^{M}$. In order to obtain the boxes with the corresponding semantics, we project the 3D boxes back to the 2D image planes $I_{t}$ and crop the corresponding image regions according to the projected box corners. The cropped images are then fed into a pre-trained CLIP model $\Theta$ [33] to obtain the semantic features $\{F_{i}\}_{i=1}^{M}$ of each box proposal as shown in Eq. 1.
$F^{i}_{t}=\Theta(P(I_{t},K_{t},\{{B}^{i}_{t}\})),$ (1)
where $K_{t}$ denotes the camera intrinsics and $P$ indicates the projection and image cropping operation. $i$ and $t$ denote the box index and frame index.
@@ 3.3 Association Module
In this section, we introduce the box association module to associate the 3D bounding boxes of the same instance across different views. The association is crucial for maintaining consistent global 3D bounding boxes since numerous bounding boxes are generated from each single view. Previous methods [44, 38] mainly exploit spatial overlap for association, which is valid for most room-level common objects like chairs and desks, but can not generalize to small objects (e.g., a TV remote on the table). The reason is that proposals for most small objects generated from different methods are usually adjacent but not spatially overlapped, which may result in the missing association and detection.
To address this challenge, we propose a novel two-stage association module: 3D Non-Maximum Suppression (NMS) for associating spatially overlapping boxes, and a 2D correspondence matching module for associating the remaining fine-grained objects’ boxes. The spatial association and correspondence association are introduced as follows.
Figure 3: Visualization of multi-view box fusion using random optimization with the IoU of the intersecting convex hulls of projected corners between the global box and single-view proposals. Three proposals of a cabinet (bottom) are fused into a global one (top).
@@ 3.3.1 Spatial Association
Concretely, given the local box proposals $B_{t}$, we add these to global bounding boxes $\mathcal{G}_{t-1}$ using the camera pose and derive temporally current global boxes $\mathcal{G}_{t}$ at frame $t$. We need to determine whether any of the global boxes correspond to each other and remove the redundant ones. Compared to the axis-aligned NMS used in previous methods [30, 26], we aim to generate scene-level oriented 3D bounding boxes. Therefore, we adopt the NMS for oriented bounding boxes. To be specific, we first sort $\mathcal{G}_{t}$ by the corresponding scores from Section 3.2. For each box $\mathcal{G}^{i}_{t}$ indexed by the descending order of scores, we calculate the IoU between $\mathcal{G}^{i}_{t}$ and the other boxes in $\mathcal{G}_{t}$. The IoU is calculated based on the convex hull of the 3D bounding boxes, which is more suitable for oriented boxes.
For online efficiency, we uniformly sample $O_{n}$ points in the convex hull of each pair of 3D bounding boxes, and only those that are likely to have overlap are considered. At first, this can be verified by checking whether the points on the box edges lie with each other. For brevity, we omit the time stamp $t$ in the following. For a pair of 3D bounding boxes $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$ ($i>j$ in the descending order of scores), the 3D convex-based IoU $\Omega_{ij}$ is defined as Eq. 2. We manage a candidate list $\Psi_{i}$ for $\mathcal{G}_{i}$. We filter $\mathcal{G}_{j}$ if $\Omega_{ij}$ is higher than the threshold $\tau_{3d}$ and $\mathcal{G}_{j}$ is then associated with $\mathcal{G}_{i}$ if the camera direction is larger than $\tau_{r}$ or the camera translation is larger than $\tau_{t}$ compared to all candidate boxes in $\Psi_{i}$.
$\Omega_{ij}=\frac{\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{G}_{i}\cup\mathcal{G}_{j})}{\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{G}_{i})+\sum_{k=1}^{O_{n}}\mathcal{C}(O_{k},\mathcal{G}_{j})},$ (2)
where $\mathcal{C}$ is the indicator function that returns 1 if the point $O_{k}\in\mathcal{G}_{i}$, otherwise 0 and $\mathcal{G}_{i}\cup\mathcal{G}_{j}$ denotes the union of convex hulls of 3D bounding box $\mathcal{G}_{i}$ and $\mathcal{G}_{j}$.